OpenStudy (anonymous):

if $12000 is invested at an interest rate of 10% per year compounded monthly. given a(t)=p(1+r over n)^nt. find the amount of the investment at the end of 3 years, and how long will the investment take to double?

OpenStudy (anonymous):

a(t)=p( 1+ (r/n) )^(nt) a(3)=12000(1+ (0.10/12) )^(3*12)

OpenStudy (anonymous):

would u be able to tell me what the final answer is because im not sure if i did the equation right

OpenStudy (anonymous):

whad did you get for the value?

OpenStudy (anonymous):

a(3)=12000(1+ (0.10/12) )^(3*12) a(3)=12000(1+ (0.10/12) )^(36) a(3)=12000(1+ 0.00833 )^(36) a(3)=12000(1.00833)^(36) a(3)=12000(1.348)

OpenStudy (anonymous):

-0.001835789

OpenStudy (anonymous):

i see what i did wrong now it makes sense

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

how long would it take the investment to double? do you know what the formula set up is for that?

OpenStudy (anonymous):

ya, so we are looking for time, or t, rather than before where we were solving for a(t). so a(t) = double 12000 or a(t) = 24000 24000=12000(1+ (0.10/12) )^(t*12)

OpenStudy (anonymous):

so the answer is t = 20.9 years correct?

OpenStudy (anonymous):

t = 6.9 :(

OpenStudy (anonymous):

24000=12000(1+ (0.10/12) )^(t*12) 24000=12000(1.00833 )^(t*12) 24000/12000 = 1.00833^(12t) 2 = 1.00833^(12t) log(2) = log( 1.00833^(12t)) log(2) = 12t*log( 1.00833) log(2) / log(1.00833) = 12t (log(2) / log(1.00833) )/12= t